give the equation for a line with a slope of 5 that passes through the point (4,1)
Answer by nerdybill(7384)

give the equation for a line with a slope of 5 that passes through the point (4,1)
.
Plug the given information into the "point-slope" form:
y - y1 = m(x - x1)
y - 1 = 5(x - 4)
y - 1 = 5x - 20
y = 5x - 21 (this is what they're looking for - in "slope-intercept" form)